The resistivity of a given wire of cross sectional area 0.7mm^2 is 4.9×10^-4mm. Calculate the resistance of a 2m lenght of the wire.
1
Expert's answer
2020-07-08T13:12:27-0400

The resistance "R" and resistivity "\rho" of the wire are connected with the following relation:


"R = \rho\dfrac{L}{A}"

where "L = 2m" is the lenght of the wire and "A = 0.7mm^2 = 0.7\cdot 10^{-6}m^2" is the cross sectional area of the wire.

Substituting the numerical values, get:

"R = 4.9\cdot 10^{-7}\Omega\cdot m\cdot \dfrac{2m}{0.7\cdot 10^{-6}m^2} = 1.4\Omega"

Answer. R = 1.4 Ohms.